# Accelerating Volumetric Medical Image Annotation via Short-Long Memory SAM 2

### v1

@@ 3.2.2 Dynamic Attention Fuser
The two output attentions are subsequently fused using a lightweight, learnable fusion module $\mathcal{F}$. It contains an optional projection layer followed by two ConvNeXt blocks [42], as shown in Fig. 1 (b).
$A_{SLM-SAM2}=\mathcal{F}(A_{short},A_{long})$ (4)
The fused attention $A_{SLM-SAM2}$ is of the same shape as $A_{SAM2}$, allowing it to be directly fed into the mask decoder to generate output predictions. All newly introduced modules adopt the same architecture as the original SAM 2 components to leverage the pretrained weights.
@@ 3.3 Annotation Propagation via SLM-SAM 2
The finetuning process of SLM-SAM 2 is described in Algorithm 1. The loss function incorporates focal loss, Dice loss, MAE loss for IoU prediction, and cross-entropy loss, with weights the same as those used in SAM 2 [12]. At inference, the central slice of the object of interest is selected as the initial annotation point, and the mask is propagated in a bi-directional manner, first toward one direction and subsequently in the opposite direction. To enhance the model’s ability to track the object in both directions, we randomly flip half of the sequence order during training.
@@ 4 Experiments
We have performed a thorough evaluation of the proposed SLM-SAM 2. In this section, we describe datasets, implementation details, and baselines, followed by results including ablation studies to verify the key designs of SLM-SAM 2.
@@ 4.1 Datasets
We evaluated SLM-SAM 2 on three publicly available datasets spanning different imaging modalities, covering anatomical structures including organs, bones, and muscles.
@@ 4.1.3 Muscle Segmentation
In all experiments, 5 volumes from each dataset were randomly chosen for training, and the remaining volumes were used for testing. All 2D slices were extracted from axial view.
@@ 4.2 Implementation Details
We followed the default settings of SAM 2 for end-to-end finetuning SLM-SAM 2, with the exception that only mask prompts were utilized, given that our study focuses on VOS tasks. The batch size was set to 1, and each input sequence was composed of 8 randomly selected consecutive 2D slices. For each segmentation task, the model was finetuned using the AdamW optimizer for 100 epochs. All models were trained on a single NVIDIA GeForce RTX A6000. The initial learning rates for the image encoder and the remaining model components were set to $3\times 10^{-6}$ and $5\times 10^{-6}$, respectively, and were updated using a cosine annealing scheduler. All 2D images were resized to a resolution of 1024 and normalized to a range of 0 to 255 at the volume level  [^2] For CT scans, intensities were clipped to the range [-1024, 2000] in Hounsfield Units (HU). The training data augmentation pipeline followed the SAM 2 implementation.
@@ 4.3 Baseline Methods
To evaluate the effectiveness of SLM-SAM 2, we compared its segmentation performance against state-of-the-art VOS models, including SAM 2 [12] and Cutie [31]. Additionally, we assessed its performance relative to leading 2D and 3D fully automated segmentation models, such as nnUNet [19], UNETR [39] and Swin UNETR [40]. We also compared our performance against unsupervised VOS methods, including VoxelMorph [45] and Sli2Vol [8]. Cutie leverages a query-based object transformer to integrate top-down object-level memory with bottom-up pixel-level features, segmenting foreground targets through the foreground-background masked attention. Cutie+, a variant of Cutie with larger memory bank that offers improved performance but with slower inference speed, is also included. We chose base model for Cutie and Cutie+, initialized with pretrained weights, and followed the default training settings from official GitHub. nnUNet is a robust, self-adapting framework based on the UNet architecture [4]. We evaluated the 2D nnUNet in two settings: the basic (B) and the exhaustive (E) setting. In the basic setting, the model was trained on $n$ fully annotated volumes and evaluated directly on the test set. In the exhaustive setting, we retrained a model individually for each test volume, using $n$ fully annotated volumes along with one annotated slice from that test volume. UNETR and SwinUNETR are state-of-the-art 3D transformer-based medical image segmentation models. UNETR employs a pure transformer encoder for volumetric representation learning, while Swin UNETR utilizes swin transformer-based blocks for enhanced feature extraction. All 3D models were trained on $n$ fully annotated volumes using the AdamW optimizer with an initial learning rate of $1\times 10^{-3}$.
@@ 4.4 Evaluation Metrics
For both 5-Volume and 1-Volume setting, we estimated model robustness using bootstrapping with 1,000 resamples. For the 1-Volume setting, we averaged performance from five models per sample. To ensure fair volume-wise evaluation, the predictions on the selected annotated slices are replaced with ground truth for nnUNet (B), UNETR and SwinUNETR.
@@ 4.5.1 Overall Performance Comparison of SLM-SAM 2
[…]
MRI-Kidney MRI-Pancreas CT-Pancreas CT-Spleen CT-Heart CT-Femur MRI-Gracilis MRI-Sartorius
Methods DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$)
@@
UNETR 0.4661 22.5008 0.2754 13.3468 0.0320 76.1877 0.1548 51.7418 0.5727 21.5100 0.5524 10.3962 0.4691 3.8774 0.3260 7.7069
SwinUNETR 0.6508 4.6814 0.3778 8.8737 0.1525 13.4887 0.3182 21.0385 0.5937 17.1132 0.5948 8.6549 0.5324 3.9244 0.4981 3.7137
III. Unsupervised VOS Models
VoxelMorph 0.3854 5.8693 0.2501 13.7858 0.4307 15.6178 0.7806 2.6862 0.5211 6.6286 0.3411 8.3026 0.5707 3.0872 0.5163 2.8654
Sli2Vol 0.8287 1.6272 0.4676 10.9631 0.4636 13.0736 0.7549 5.3969 0.6243 6.3413 0.3042 9.1010 0.5214 3.7066 0.4352 4.0077
IV. Semi-supervised VOS Models
Cutie 0.8976 1.9001 0.5340 6.6390 0.6806 8.3856 0.9629 0.5550 0.7855 6.2811 0.6303 16.7382 0.7081 2.1894 0.6813 1.5448
Cutie+ 0.8997 1.6683 0.5572 6.3690 0.6628 7.7925 0.9574 0.5197 0.7031 12.2835 0.6285 17.0783 0.7209 2.1434 0.6839 1.5357
SAM 2 0.8668 5.9103 0.2341 30.4152 0.6381 9.1385 0.9633 0.4508 0.6925 15.6042 0.6267 52.3565 0.7926 1.4828 0.7045 1.4530
SLM-SAM 2 0.9368 1.3643 0.6294 8.4185 0.7255 4.0753 0.9607 0.4767 0.7655 8.2143 0.9643 0.4298 0.8543 0.7276 0.8118 0.7524
@@ IV. Semi-supervised VOS Models
The average performance of SLM-SAM 2 with other baselines is presented in Fig. 2. Overall, we observed that SLM-SAM 2 consistently outperforms the competing methods by a significant margin, achieving improvements of over 0.07 in DSC and 2.99 in ASSD at least. In particular, while SAM 2 achieves a DSC of 0.69, SLM-SAM 2 attains a DSC of 0.83, thereby demonstrating the benefits of incorporating both short-term and long-term memory banks in mask propagation.
VOS methods generally show superior performance in slice propagation tasks while maintaining comparable computational costs. One explanation is both Cutie and SAM 2 are initialized with pretrained weights, leading to outperformance when finetuned on small datasets. Notably, Cutie and Cutie+ slightly outperformed SAM 2, achieving DSC of 0.73 and 0.72, respectively. This is reasonable as Cutie is found to be one of the most competitive methods relative to SAM 2 [12].
The runner-up method, nnUNet (E), requires retraining for each new volume, which leads to a considerable increase in time and computational resources. However, excluding the annotated slice in the training (i.e., nnUNet (B)) results in a significant performance degradation (0.19 reduction in DSC).
The 3D models, UNETR and SwinUNETR, show suboptimal performance with DSC values of 0.38 and 0.48, respectively. One reason could be they do not leverage annotated slice information during either training or inference stages.
Moreover, unsupervised methods, VoxelMorph and Sli2Vol, exhibit lower DSC performance compared to the VOS-based methods. This underperformance can be attributed to the training on unlabeled data, which focuses on learning general information transformation between slices and may not directly learn task-specific features for accurate segmentation.
@@ 4.5.2 Performance Comparison of Each Dataset/Task
The quantitative results (DSC and ASSD) of SLM-SAM 2 and all competing methods across the datasets/tasks are presented in Table 1 and Fig. 3. Overall, SLM-SAM 2 outperforms the other baselines in terms of DSC on most datasets. For the CT-Spleen tasks, all VOS methods achieve a DSC of around 0.96, performing comparable to each other. On CT-Heart, SLM-SAM 2 (DSC: 0.7655) underperforms nnUNet (E) (DSC: 0.8013) and slightly underperforms Cutie (DSC: 0.7855). However, nnUNet (E) requires retraining for each new volumes, limiting its scalability. It is notably that SLM-SAM 2 consistently outperforms SAM 2 across organs, bones, and muscles, demonstrating its improved robustness in mask propagation tasks. Moreover, among 2D method, nnUNet (E) consistently outperform nnUNet (B) for all datasets. This is expected as nnUNet (E) leverages an additional annotated slice from each test volume. For the 3D methods, SwinUNETR demonstrates superior performance compared to UNETR on all datasets. For unsupervised VOS methods, VoxelMorph outperforms Sli2Vol on certain datasets, such as CT-Femur and MRI muscle tasks while Sli2Vol achives better performance than VoxelMorph on both MRI and CT organ tasks.
@@
For target-absent slices (i.e., slices where target is absent), SAM 2 and several other VOS methods tends to over-propagate masks of segmentation into regions where the target is no longer present, particularly in adjacent anatomical structures with similar intensity. For instance, in MRI-Kidney, Cutie, Cutie+ and SAM 2 sometimes over-propagate the kidney masks onto the spleen. Similarly, in CT-Femur, other VOS methods may segment the hip bones as the femurs, likely due to their similar appearance and spatial overlap in axial 2D slices. These challenges are not limited to VOS methods. Competing 2D and 3D segmentation models, such as nnUNet (E), UNETR and SwinUNETR, can also mis-segment adjacent anatomical structures. For example, these models may incorrectly segment the spleen in MRI-Kidney tasks. However, experimental results demonstrate that SLM-SAM 2, which incorporates dynamically learned short-term and long-term memory mechanisms, can effectively alleviate this issue (Fig. 4, Fig. 5 (a) and Fig. 6), leading to more accurate and robust models for slice propagation.
@@
We further investigate the benefits of integrating both short-term and long-term memory in SLM-SAM 2, compared to SAM 2. To do this, we compute the 2D slice-wise DSC on target-present slices and target-absent slices, as shown in Fig. 5 (a).
@@
For most datasets except two MRI muscles datasets, SLM-SAM 2 and SAM 2 achieve comparable performance on target-present slices. However, for target-absent slices, SLM-SAM 2 demonstrates significant outperformance compared to SAM 2, with particularly notable improvements observed on MRI-Kidney, MRI-Pancreas, CT-Pancreas and CT-Femur.
Fig. 5 (b)-(e) presents more direct and illustrative results. Specifically, Fig. 5 (b) depicts the 2D DSC over slice index for an example from MRI-Kidney dataset. Figs. 5 (c), (d) and (e) provide qualitative comparison of the segmentation performance between SLM-SAM 2 versus SAM 2 on the selected representative slices in the MRI-Kidney example. Firstly, it can be seen that the 2D DSC on target-present slices is similar between SLM-SAM 2 and SAM 2, as shown in Fig. 5 (c). However, once the kidney is no longer present (i.e., beyond the gray panel range), the 2D DSC of SAM 2 drops to zero, whereas SLM-SAM 2 maintains a 2D DSC of 1. As illustrated in Fig. 5 (d), this is because SAM 2 over-propagates kidney mask into the spleen region. A potential explanation is the similar intensity between the spleen and kidney, which may obscure their boundaries. Furthermore, as depicted in Fig. 5 (e), SAM 2 stops propagating the mask after the spleen disappears from the images. This may be because the adjacent lung tissue has a much lower intensity (i.e., darker) in MRI, making the boundary more distinct.
On target-present slices, SLM-SAM 2 achieves an overall better performance compared to SAM 2 in MRI muscle datasets, while maintaining comparable performance across other tasks. By incorporating a dynamic learning mechanism that integrates both short-term and long-term memory, SLM-SAM 2 exhibits improved robustness at target boundaries, significantly mitigating over-propagation problem. We argue that such vague boundary situations are common in medical imaging, like unclear boundaries often occur between the femur and hip bones in CT scans, as shown in Fig. 4. These challenges may result from factors such as intensity similarities between the target and neighboring structures, close proximity or direct contact between anatomical structures, and limitations in scan resolution (i.e., slice thickness or spacing between slices). Therefore, the ability of SLM-SAM 2 to better handle boundary ambiguity makes it a promising approach for medical image segmentation.
@@
MRI-Kidney MRI-Pancreas CT-Pancreas CT-Spleen CT-Heart CT-Femur MRI-Gracilis MRI-Sartorius
Methods $\mathcal{M}_{1}$ $\mathcal{M}_{2}$ DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$)
$\mathcal{M}_{O}$ (No FT) $\mathcal{M}_{O}$ 0.6428 0.3358 0.1345 0.2858 0.4271 0.4359 0.5662 0.3483
$\mathcal{M}_{O}$ $\mathcal{M}_{O}$ 0.8668 0.2341 0.6381 0.9633 0.6925 0.6267 0.7926 0.7045
$\mathcal{M}_{R}=7$ $\mathcal{M}_{R}=7$ 0.9088 0.3009 0.6552 0.9590 0.6650 0.9232 0.7043 0.4989
$\mathcal{M}_{R}=1$ $\mathcal{M}_{R}=1$ 0.9018 0.4866 0.6637 0.9494 0.7138 0.9351 0.7900 0.7882
$\mathcal{M}_{O}+\mathcal{M}_{R}=7$ $\mathcal{M}_{O}$ $\mathcal{M}_{R}=7$ 0.8916 0.3765 0.6690 0.9621 0.7177 0.9559 0.7765 0.6836
$\mathcal{M}_{O}+\mathcal{M}_{R}=1$ $\mathcal{M}_{O}$ $\mathcal{M}_{R}=1$ 0.9368 0.6294 0.7255 0.9607 0.7655 0.9643 0.8543 0.8118

### v2

@@ 3.2.2 Dynamic Attention Fuser
The two output attentions are of the same shape $(B,C,H,W)$ with $C=256,H=W=64$ and are subsequently fused using a lightweight, learnable fusion module $\mathcal{F}$. The fusion module first performs element-wise summation of $A_{short}$ and $A_{long}$, followed by an optional projection layer and two ConvNeXt blocks [51], as shown in Fig. 1 (b). Each ConvNeXt begins with a depthwise $7\times 7$ convolution followed by LayerNorm. The normalized output is then permuted to $(B,H,W,C)$ and passed through a two-layer pointwise MLP that expands channels from $C$ to $4C$, applies GELU, then projects back to $C$. Finally, the output is permuted back to $(B,C,H,W)$ and combined with the original input via a residual connection. Simply, $A_{SLM-SAM2}$ can be computed as:
$(A_{SLM-SAM2})_{t}=\mathcal{F}((A_{short})_{t},(A_{long})_{t})$ (8)
Thus, the fused attention $A_{SLM-SAM2}$ is of the same shape as $A_{SAM2}$, allowing it to be directly fed into the mask decoder to generate output predictions. All newly introduced modules adopt the same architecture as the original SAM 2 components to leverage the pretrained weights.
@@ 3.3 Annotation Propagation via SLM-SAM 2
The finetuning process of SLM-SAM 2 is described in Algorithm 1. The loss function incorporates focal loss, Dice loss, MAE loss for IoU prediction, and cross-entropy loss, with weights the same as those used in SAM 2 [12]. At inference, the middle slice of the object of interest is selected as the initial annotation point, and the mask is propagated in a bi-directional manner, first toward one direction and subsequently in the opposite direction. To enhance the model’s ability to track the object in both directions, we randomly flip half of the sequence order during training.
@@ 4 Experiments
We have performed a thorough evaluation of the proposed SLM-SAM 2. In this section, we describe datasets, implementation details, and baselines, followed by results including ablation studies to verify the key designs of SLM-SAM 2, propagation error analysis, and quantitative efficiency experiments.
@@ 4.1 Datasets
We evaluated SLM-SAM 2 on four publicly available datasets spanning different imaging modalities, covering anatomical structures including organs, bones, and muscles.
@@ 4.1.3 Muscle Segmentation
4.1.4 Ultrasound Segmentation
: The JNU-IFM (Intelligent Fetal Monitoring Lab of Jinan University) dataset [55] was used to evaluate fetal head segmentation in ultrasound videos (US-Head), with 25 randomly selected sequences.
In all experiments, 5 volumes from each dataset were randomly chosen for training, and the remaining volumes were used for testing. All 2D slices (except US-Head) were extracted from axial view.
@@ 4.2 Implementation Details
We followed the default settings of SAM 2 for end-to-end finetuning SLM-SAM 2, with the exception that only mask prompts were utilized, given that our study focuses on VOS tasks. All components were trainable during finetuning and all parameters were optimized jointly. The batch size was set to 1, and each input sequence was composed of 8 randomly selected consecutive 2D slices. For each segmentation task, the model was finetuned using the AdamW optimizer for 100 epochs. Due to the limited number of training samples, we evaluated using the model checkpoint from the final training epoch, in a setting similar to [56, 46]. All models were trained on a single NVIDIA GeForce RTX A6000. The initial learning rates for the image encoder and the remaining model components were set to $3\times 10^{-6}$ and $5\times 10^{-6}$, respectively, and were updated using a cosine annealing scheduler. All 2D images were resized to a resolution of 1024 and normalized to a range of 0 to 255 at the volume level  [^3] For CT scans, intensities were clipped to the range [-1024, 2000] in Hounsfield Units (HU). The training data augmentation pipeline followed the SAM 2 implementation.
@@ 4.3 Baseline Methods
To evaluate the effectiveness of SLM-SAM 2, we compared its segmentation performance against state-of-the-art VOS models, including SAM 2 [12], Cutie [41] and iSegFormer [38]. Additionally, we assessed its performance relative to leading 2D and 3D fully automated segmentation models, such as nnUNet [19], Medical SAM Adapter (MSA) [13], UNETR [47] and Swin UNETR [48]. Furthermore, we compared with ConvLSTM-based and BiConvLSTM-based methods, such as LSTM-MM-UNet [57] and BCDUNet [58]. We also compared our performance against unsupervised VOS methods, including VoxelMorph [59] and Sli2Vol [8]. VoxelMorph and Sli2Vol are image-registration methods that learn dense correspondences between adjacent slices to propagate manual annotations across an entire volume. Cutie leverages a query-based object transformer to integrate top-down object-level memory with bottom-up pixel-level features, segmenting foreground targets through the foreground-background masked attention. Cutie+, a variant of Cutie with a larger memory bank that offers improved performance but with slower inference speed, is also included. We chose the base model for Cutie and Cutie+, initialized with pretrained weights, and followed the default training settings from official GitHub. iSegFormer integrates a 2D interactive segmentation network with a label-propagation module. In our experiments, we evaluated its propagation component with pretrained weights initialization starting from a single annotated mask. nnUNet is a robust, self-adapting framework based on the UNet architecture [5]. We evaluated the 2D nnUNet in two settings: the basic (B) and the exhaustive (E) setting. In the basic setting, the model was trained on $n$ fully annotated volumes and evaluated directly on the test set. In the exhaustive setting, we retrained a model individually for each test volume, using $n$ fully annotated volumes along with one annotated slice from that test volume. LSTM-MM-UNet uses ConvLSTM to enforce inter-slice continuity, while BCDUNet replaces the standard skip-connection in UNet with a BiConvLSTM module. We followed the official GitHub for implementation. UNETR and SwinUNETR are state-of-the-art 3D transformer-based medical image segmentation models. UNETR employs a pure transformer encoder for volumetric representation learning, while Swin UNETR utilizes swin transformer-based blocks for enhanced feature extraction. All 3D models were trained on $n$ fully annotated volumes using the AdamW optimizer with an initial learning rate of $1\times 10^{-3}$. For fair comparison, all propagation methods propagated labels bidirectionally, used the same dataset splits and the same manual masks as the initial prompts at inference time. In addition, for automatic segmentation methods, the predictions on the selected annotated slices were replaced with the ground truth during evaluation.
@@ 4.4 Evaluation Metrics
For both 5-Volume and 1-Volume settings, we estimated model robustness using bootstrapping with 1,000 resamples. For the 1-Volume setting, we averaged performance from five models per sample.
@@ 4.5.1 Overall Performance Comparison of SLM-SAM 2
Table 1: Quantitative Comparison (DSC and ASSD) for each dataset under the 5-Volume settings. Best results are in bold; runner-up is underlined. $\dagger$}: method requires retraining for each test volume.
MRI-Kidney MRI-Pancreas CT-Pancreas CT-Spleen CT-Heart CT-Femur MRI-Gracilis MRI-Sartorius US-Head
Methods DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$) DSC ($\uparrow$) ASSD ($\downarrow$)
@@
UNETR 0.4661 22.5008 0.2754 13.3468 0.0320 76.1877 0.1548 51.7418 0.5727 21.5100 0.5524 10.3962 0.4691 3.8774 0.3260 7.7069 0.5376 18.6205
SwinUNETR 0.6508 4.6814 0.3778 8.8737 0.1525 13.4887 0.3182 21.0385 0.5937 17.1132 0.5948 8.6549 0.5324 3.9244 0.4981 3.7137 0.5796 14.2872
III. ConvLSTM/BiConvLSTM-based Models
LSTM-$M$-UNet 0.4483 18.7278 0.0741 18.4087 0.1178 21.6110 0.3521 14.5100 0.1877 46.1875 0.4406 31.9848 0.1423 54.3272 0.0935 51.4670 0.5311 30.1221
BCDUNet 0.4253 4.4763 0.0651 21.7007 0.1100 26.6618 0.1191 7.2795 0.4579 25.5236 0.0539 15.8434 0.2282 4.4915 0.2674 3.3482 0.4005 10.8755
IV. Unsupervised VOS Models
VoxelMorph 0.3854 5.8693 0.2501 13.7858 0.4307 15.6178 0.7806 2.6862 0.5211 6.6286 0.3411 8.3026 0.5707 3.0872 0.5163 2.8654 0.4984 10.8782
Sli2Vol 0.8287 1.6272 0.4676 10.9631 0.4636 13.0736 0.7549 5.3969 0.6243 6.3413 0.3042 9.1010 0.5214 3.7066 0.4352 4.0077 0.4128 15.3546
V. Semi-supervised VOS Models
iSegFormer 0.8067 8.9592 0.4358 17.7530 0.5016 14.4014 0.8605 7.8476 0.6496 13.9058 0.5769 47.0773 0.8156 1.7567 0.7347 4.5420 0.5239 51.1868
Cutie 0.8976 1.9001 0.5340 6.6390 0.6806 8.3856 0.9629 0.5550 0.7855 6.2811 0.6303 16.7382 0.7081 2.1894 0.6813 1.5448 0.8500 3.1714
Cutie+ 0.8997 1.6683 0.5572 6.3690 0.6628 7.7925 0.9574 0.5197 0.7031 12.2835 0.6285 17.0783 0.7209 2.1434 0.6839 1.5357 0.8501 3.1133
SAM 2 0.8668 5.9103 0.2341 30.4152 0.6381 9.1385 0.9633 0.4508 0.6925 15.6042 0.6267 52.3565 0.7926 1.4828 0.7045 1.4530 0.8635 2.3445
SLM-SAM 2 0.9368 1.3643 0.6294 8.4185 0.7255 4.0753 0.9607 0.4767 0.7655 8.2143 0.9643 0.4298 0.8543 0.7276 0.8118 0.7524 0.8764 2.4207
@@ V. Semi-supervised VOS Models
The average performance of SLM-SAM 2 with other baselines is presented in Fig. 2. Overall, we observed that SLM-SAM 2 consistently outperforms the competing methods by a significant margin, achieving improvements of over 0.07 in DSC and 2.77 in ASSD at least. In particular, while SAM 2 achieves a DSC of 0.70, SLM-SAM 2 attains a DSC of 0.84, thereby demonstrating the benefits of incorporating both short-term and long-term memory banks in mask propagation.
VOS methods generally show superior performance in slice propagation tasks while maintaining comparable computational costs. One explanation is iSegFormer, Cutie and SAM 2 are initialized with pretrained weights, leading to outperformance when finetuned on small datasets. Notably, Cutie and Cutie+ slightly outperformed SAM 2, achieving DSC of 0.74 and 0.73, respectively. This is reasonable as Cutie is found to be one of the most competitive methods relative to SAM 2 [12].
The runner-up method, nnUNet (E), requires retraining for each new volume, which leads to a considerable increase in time and computational resources. However, excluding the annotated slice in the training (i.e., nnUNet (B)) results in a significant performance degradation (0.19 reduction in DSC). Moreover, MSA, which leverages the SAM’s pretrained encoder, exhibits superior performance than nnUNet (B).
The 3D models, UNETR and SwinUNETR, show suboptimal performance with DSC values of 0.39 and 0.49, respectively. One reason could be they do not leverage annotated slice information during either the training or inference stages.
Moreover, unsupervised methods, VoxelMorph and Sli2Vol, exhibit lower DSC performance compared to the VOS-based methods. This underperformance can be attributed to the training on unlabeled data, which focuses on learning general information transformation between slices and may not directly learn task-specific features for accurate segmentation. Additionally, ConvLSTM-based methods, LSTM-MM-UNet and BCDUNet, built on UNet backbones without pretrained encoders, yield the lowest segmentation accuracies compared to other evaluated approaches.
@@ 4.5.2 Performance Comparison of Each Dataset/Task
The quantitative results (DSC and ASSD) of SLM-SAM 2 and all competing methods across the datasets/tasks are presented in Table 1 and Fig. 3. Overall, SLM-SAM 2 outperforms the other baselines in terms of DSC on most datasets. For the CT-Spleen tasks, all VOS methods achieve a DSC of around 0.96, performing comparable to each other. On CT-Heart, SLM-SAM 2 (DSC: 0.7655) underperforms nnUNet (E) (DSC: 0.8013) and slightly underperforms Cutie (DSC: 0.7855). However, nnUNet (E) requires retraining for each new volume, limiting its scalability. It is notably that SLM-SAM 2 consistently outperforms SAM 2 across organs, bones, muscles, and fetal head, demonstrating its improved robustness in mask propagation tasks. Moreover, among 2D method, nnUNet (E) consistently outperform nnUNet (B) for all datasets. This is expected as nnUNet (E) leverages an additional annotated slice from each test volume. For the 3D methods, SwinUNETR demonstrates superior performance compared to UNETR on all datasets. For unsupervised VOS methods, VoxelMorph outperforms Sli2Vol on certain datasets, such as CT-Femur, MRI muscle tasks and US-Head while Sli2Vol achieves better performance than VoxelMorph on both MRI and CT organ tasks. For ConvLSTM-based methods, BCDUNet outperforms LSTM-MM-UNet on CT-Heart and MRI muscle tasks, while LSTM-MM-UNet achieves superior performance on the others.
@@
For target-absent slices (i.e., slices where the target is absent), SAM 2 and several other VOS methods tend to over-propagate masks of segmentation into regions where the target is no longer present, particularly in adjacent anatomical structures with similar intensity. For instance, in MRI-Kidney, iSegFormer, Cutie, Cutie+ and SAM 2 sometimes over-propagate the kidney masks onto the spleen. Similarly, in CT-Femur, other VOS methods may segment the hip bones as the femurs, likely due to their similar appearance and spatial overlap in axial 2D slices. These challenges are not limited to VOS methods. Competing 2D and 3D segmentation models, such as nnUNet (E), UNETR and SwinUNETR, can also mis-segment adjacent anatomical structures. For example, these models may incorrectly segment the spleen in MRI-Kidney tasks. However, experimental results demonstrate that SLM-SAM 2, which incorporates dynamically learned short-term and long-term memory mechanisms, can effectively alleviate this issue (Fig. 4, Fig. 5 (a) and Fig. 6), leading to more accurate and robust models for slice propagation.
@@
We further investigated the benefits of integrating both short-term and long-term memory in SLM-SAM 2, compared to SAM 2. To do this, we computed the 2D slice-wise DSC on target-present slices and target-absent slices, as shown in Fig. 5 (a).
@@
For most datasets except two MRI muscles datasets, SLM-SAM 2 and SAM 2 achieve comparable performance on target-present slices. However, for target-absent slices, SLM-SAM 2 demonstrates significant outperformance compared to SAM 2, with particularly notable improvements observed on MRI-Kidney, MRI-Pancreas, CT-Pancreas and CT-Femur. In addition, SLM-SAM 2 exhibits superior performance on both target-present and target-absent slices on US-Head.
Fig. 5 (b)-(e) presents more direct and illustrative results. Specifically, Fig. 5 (b) depicts the 2D DSC over the slice index for an example from MRI-Kidney dataset. Fig. 5 (c), (d) and (e) provide a qualitative comparison of the segmentation performance between SLM-SAM 2 versus SAM 2 on the selected representative slices in the MRI-Kidney example. Firstly, it can be seen that the 2D DSC on target-present slices is similar between SLM-SAM 2 and SAM 2, as shown in Fig. 5 (c). However, once the kidney is no longer present (i.e., beyond the gray panel range), the 2D DSC of SAM 2 drops to zero, whereas SLM-SAM 2 maintains a 2D DSC of 1. As illustrated in Fig. 5 (d), this is because SAM 2 over-propagates the kidney mask into the spleen region. A potential explanation is the similar intensity between the spleen and kidney, which may obscure their boundaries. Furthermore, as depicted in Fig. 5 (e), SAM 2 stops propagating the mask after the spleen disappears from the images. This may be because the adjacent lung tissue has a much lower intensity (i.e., darker) in MRI, making the boundary more distinct.
On target-present slices, SLM-SAM 2 achieves an overall better performance compared to SAM 2 in MRI muscle datasets and US-Head, while maintaining comparable performance across other tasks. By incorporating a dynamic learning mechanism that integrates both short-term and long-term memory, SLM-SAM 2 exhibits improved robustness at target boundaries, significantly mitigating over-propagation problem. We argue that such vague boundary situations are common in medical imaging, like unclear boundaries often occur between the femur and hip bones in CT scans, as shown in Fig. 4. These challenges may result from factors such as intensity similarities between the target and neighboring structures, close proximity or direct contact between anatomical structures, and limitations in scan resolution (i.e., slice thickness or spacing between slices). Therefore, the ability of SLM-SAM 2 to better handle boundary ambiguity makes it a promising approach for medical image segmentation.
@@
MRI-Kidney MRI-Pancreas CT-Pancreas CT-Spleen CT-Heart CT-Femur MRI-Gracilis MRI-Sartorius US-Head
Methods $\mathcal{M}_{1}$ $\mathcal{M}_{2}$ DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$) DSC ($\uparrow$)
$\mathcal{M}_{O}$ (No FT) $\mathcal{M}_{O}$ 0.6428 0.3358 0.1345 0.2858 0.4271 0.4359 0.5662 0.3483 0.8233
$\mathcal{M}_{O}$ $\mathcal{M}_{O}$ 0.8668 0.2341 0.6381 0.9633 0.6925 0.6267 0.7926 0.7045 0.8635
$\mathcal{M}_{R}=7$ $\mathcal{M}_{R}=7$ 0.9088 0.3009 0.6552 0.9590 0.6650 0.9232 0.7043 0.4989 0.8270
$\mathcal{M}_{R}=1$ $\mathcal{M}_{R}=1$ 0.9018 0.4866 0.6637 0.9494 0.7138 0.9351 0.7900 0.7882 0.8261
$\mathcal{M}_{O}+(\mathcal{M}_{R}=7)$ $\mathcal{M}_{O}$ $\mathcal{M}_{R}=7$ 0.8916 0.3765 0.6690 0.9621 0.7177 0.9559 0.7765 0.6836 0.8538
$\mathcal{M}_{O}+(\mathcal{M}_{R}=1)$ $\mathcal{M}_{O}$ $\mathcal{M}_{R}=1$ 0.9368 0.6294 0.7255 0.9607 0.7655 0.9643 0.8543 0.8118 0.8764